# An Introduction to Reinforcement Learning: Fundamental Concepts and Practical Applications

### v1

@@ 4.3 Policy Evaluation (Prediction)
Policy evaluation, also known as prediction, involves computing the state-value function $v^{\pi}$ for a given policy $\pi$. This process assesses the expected return when following policy $\pi$ from each state. The state-value function $v_{\pi}(s)$ is defined as the expected return starting from state $s$ and following policy $\pi$:
$v_{\pi}(s)\equiv\mathbb{E}_{\pi}[R_{t+1}+\gamma G_{t+1}\mid S_{t}=s]$ (28)
@@
$v_{\pi}(s)=\mathbb{E}_{\pi}[R_{t+1}+\gamma v_{\pi}(S_{t+1})\mid S_{t}=s]=\sum_{a}\pi(a\mid s)\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)[r+\gamma v_{\pi}(s^{\prime})]$ (29)
In these equations, $\pi(a\mid s)$ denotes the probability of taking action $a$ in state $s$ under policy $\pi$. The existence and uniqueness of $v^{\pi}$ are guaranteed if $\gamma<1$ or if all states eventually terminate under $\pi$. Dynamic Programming (DP) algorithm updates are termed ”expected updates” because they rely on the expectation over all potential next states, rather than just a sample [sutton2018reinforcement].
@@ 4.4 Policy Improvement
The purpose of calculating the value function for a policy is to identify improved policies. Assume $v^{\pi}$ for a deterministic policy $\pi$. For a state $s$, should we alter the policy to select action $a\neq\pi(s)$? We know the effectiveness of adhering to the existing policy from state $s$ ($v^{\pi}(s)$), but would transition to a new policy yield a superior outcome? We can answer this by selecting action $a$ in $s$ and then following $\pi$:
To determine if a policy can be improved, we compare the value of taking a different action $a$ in state $s$ with the current policy. This is done using the action value function $q_{\pi}(s,a)$:
$q_{\pi}(s,a)\equiv\mathbb{E}[R_{t+1}+\gamma v_{\pi}(S_{t+1})\mid S_{t}=s,A_{t}=a]=\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)[r+\gamma v_{\pi}(s^{\prime})]$ (30)
The key criterion is whether this value exceeds $v^{\pi}(s)$. If $q_{\pi}(s,a)>v^{\pi}(s)$, consistently choosing action $a$ in $s$ is more advantageous than following $\pi$, leading to an improved policy $\pi^{\prime}$.
@@ 4.5 Policy Improvement Theorem
The policy improvement theorem states that if $q_{\pi}(s,\pi^{\prime}(s))\geq v_{\pi}(s)$ for all states $s$, then the new policy $\pi^{\prime}$ is at least as good as the original policy $\pi$. Formally, it is expressed as follows:
Let $\pi$ and $\pi^{\prime}$ be deterministic policies such that for all $s\in S$:
@@
If $\pi^{\prime}$ achieves greater or equal expected return from all states $s\in S$:
@@
If there is strict inequality at any state, $\pi^{\prime}$ is superior to $\pi$. Extend this to all states and actions, selecting the action that maximizes $q_{\pi}(s,a)$:
The new policy $\pi^{\prime}$ is obtained by selecting the action that maximizes the action value function $q_{\pi}(s,a)$:
$\pi^{\prime}(s)\equiv\arg\max_{a}q_{\pi}(s,a)=\arg\max_{a}\mathbb{E}[R_{t+1}+\gamma v_{\pi}(S_{t+1})\mid S_{t}=s,A_{t}=a]=\arg\max_{a}\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)[r+\gamma v_{\pi}(s^{\prime})]$ (33)
Policy improvement creates a new policy that enhances an initial policy by adopting a greedy approach based on the value function.
Assume $\pi^{\prime}$ is equally effective as $\pi$ but not superior. Then $v_{\pi}=v_{\pi^{\prime}}$, ensuring for all states $s\in S$:
The relationship between the optimal state value function $v^{*}(s)$ and the optimal action value function $q^{*}(s,a)$ is given by the following equation:
$v_{\pi^{\prime}}(s)=\max_{a}\mathbb{E}[R_{t+1}+\gamma v_{\pi^{\prime}}(S_{t+1})\mid S_{t}=s,A_{t}=a]=\max_{a}\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)[r+\gamma v_{\pi^{\prime}}(s^{\prime})]$ (34)
This is the Bellman optimality equation, implying $v_{\pi^{\prime}}=v^{*}$ and both $\pi$ and $\pi$ are optimal policies. Policy improvement yields a superior policy unless the initial policy is already optimal.
This concept extends to stochastic policies. Stochastic policies introduce a set of probabilities for actions, with the action most aligned with the greedy policy assigned the highest probability.
@@ 4.6 Policy Iteration
After enhancing a policy $\pi$ using $v_{\pi}$ to derive an improved policy $\pi^{\prime}$, compute $v_{\pi^{\prime}}$ and further refine it to obtain a superior policy $\pi^{\prime\prime}$. This process generates a sequence of improving policies and corresponding value functions:
The process of policy iteration involves alternating between policy evaluation and policy improvement to obtain a sequence of improving policies and value functions:
@@
Each policy in this sequence is a marked improvement over its predecessor unless the preceding one is already optimal. Given a finite MDP, this iterative process converges to an optimal policy and value function in a finite number of iterations. This method is called policy iteration.
5 Value Iteration
@@
One limitation of policy iteration is that each iteration requires policy evaluation, often necessitating multiple passes through the entire state set [bertsekas2011approximate]. To address this, policy evaluation can be abbreviated without losing convergence guarantees. This method, known as value iteration, terminates policy evaluation after a single sweep. It combines policy improvement with a truncated form of policy evaluation. Value iteration merges one pass of policy evaluation with policy improvement in each iteration, ensuring convergence to an optimal policy for discounted finite MDPs [lutter2021value].
The update rule for value iteration is given by:
$v_{k+1}(s)\equiv\max_{a}\mathbb{E}[R_{t+1}+\gamma v_{k}(S_{t+1})\mid S_{t}=s,A_{t}=a]=\max_{a}\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)[r+\gamma v_{k}(s^{\prime})]$ (36)
Value iteration combines one sweep of policy evaluation and policy improvement in each iteration. It converges to an optimal policy for discounted finite MDPs [lutter2021value].
Policy iteration involves two processes: policy evaluation aligns the value function with the current policy, and policy improvement makes the policy greedier based on the value function. These processes iteratively reinforce each other until an optimal policy is obtained.
@@
In value iteration, the key advantage is its efficiency, as it reduces the computational burden by merging policy evaluation and improvement into a single update step. This method is particularly useful for large state spaces where full policy evaluation at each step of policy iteration is computationally prohibitive [sutton2018reinforcement]. Additionally, value iteration can be implemented using a synchronous update approach, where all state values are updated simultaneously, or an asynchronous update approach, where state values are updated one at a time, potentially allowing for faster convergence in practice [bertsekas2011approximate].
Another notable aspect of value iteration is its robustness to initial conditions. Starting from an arbitrary value function, value iteration iteratively refines the value estimates until convergence, making it a reliable method for finding optimal policies even when the initial policy is far from optimal [bertsekas2012dynamic].
Furthermore, value iteration provides a foundation for more advanced algorithms, such as Q-learning and other reinforcement learning techniques, by illustrating the principle of bootstrapping, where the value of a state is updated based on the estimated values of successor states. This principle is central to many modern reinforcement learning algorithms that seek to balance exploration and exploitation in dynamic and uncertain environments [mnih2015human].
6 Terminology
@@
Understanding the various methodologies and concepts within RL is essential for the effective design and implementation of RL algorithms. Methods in RL can be classified as either off-policy or on-policy, as well as model-free and model-based. These categories offer different approaches and techniques for learning from interactions with the environment.
6.1 Model-Free Methods
Model-free methods determine the optimal policy or value function directly without constructing a model of the environment. There is no requirement for them to know transition probabilities and rewards, as they learn entirely from observed states, actions, and rewards. Compared with model-based methods, model-free methods are simpler to implement, relying on experience-based learning. There are two primary types:
@@ 6.1 Model-Free Methods
Value-based methods focus on learning the action-value function to derive an optimal policy. For instance, Q-learning (discussed in section 8) is an off-policy algorithm that learns the value of the optimal policy independently of the agent’s actions by using a max operator in its update rule. SARSA (also discussed in section 8), on the other hand, is an on-policy algorithm that updates its Q-values based on the actions actually taken by the policy. Both methods update their action-value estimates based on the Bellman equation until convergence. In contrast, policy-based methods, like REINFORCE (discussed in section 8), work by directly learning the policy without explicitly learning a value function. These methods adjust the policy parameters directly by following the gradient of the expected reward. This approach is particularly useful in environments with high-dimensional action spaces where value-based methods may not be effective. Policy-based methods are also capable of handling stochastic policies, providing a natural framework for dealing with uncertainty in action selection.
In addition to these primary types, there are also hybrid approaches that combine value-based and policy-based methods, such as Actor-Critic algorithms (which will be discussed in section 8). These methods consist of two components: an actor that updates the policy parameters in a direction suggested by the critic, and a critic that evaluates the action-value function. Combining both types of learning is intended to provide more stable and efficient learning [grondman2012survey].
The best way to understand hybrid approaches is to imagine a robot vacuum cleaner navigating a living room and cleaning it efficiently. It is imperative that the robot decides the best course of action in order to cover the entire area while avoiding obstacles and maximizing its battery life.
As part of the value-based component, the robot estimates the value of being in each location in the living room based on a value-based method. The robot learns a state-value function that indicates how much dirt should be removed if it starts at a specific location and follows a specific strategy. Through this component, the robot understands the long-term benefits of being in different locations. Simultaneously, the robot uses a policy-based method to determine the right action to take (for example, moving forward, turning left, turning right). As a result of the policy, the robot adjusts parameters to improve its decision-making process. For example, if moving forward usually results in more dirt being cleaned, the robot is more likely to choose this action in the future when it encounters similar circumstances.
When these two components are combined, the robot is able to navigate and clean the living room more effectively. A value-based approach provides a broader understanding of which areas are most valuable to visit, while a policy-based approach focuses on making the best decisions for the present based on the current situation. The hybrid approach ensures that the robot not only plans its long-term strategy effectively, but also reacts appropriately to immediate circumstances, resulting in an overall more efficient cleaning process.
@@
Another significant advancement in model-free methods is the development of Deep RL (DRL) By integrating deep neural networks with traditional RL algorithms, methods such as Deep Q-Networks (DQN) [mnih2013playing] and Proximal Policy Optimization (PPO) [schulman2017proximal] have achieved remarkable success in complex, high-dimensional environments, including games and robotic control tasks. The advancement of these technologies has opened up new possibilities for the application of RL to real-world problems, enabling the demonstration of robust performance in domains which were previously intractable. It is beyond the scope of this paper to discuss these algorithms, and wee refer you to [li2017deep, arulkumaran2017brief, franccois2018introduction, li2023deep] to understand DRL deeply and effectively.
6.2 Model-Based Methods
It is possible to predict the outcomes of actions using model-based methods, which facilitate strategic planning and decision-making. The use of these methods enhances learning efficiency by providing opportunities for virtual experimentation, despite the complexity of developing and refining accurate models [luo2024survey].
Autonomous driving systems are an example of how model-based methods can be applied in the real world. As autonomous vehicles navigate in dynamic environments, obstacle avoidance, and optimal routing must be made in real time.
@@ 6.2 Model-Based Methods
Autonomous vehicles create detailed models of their environment. These models include static elements, such as roads and buildings, as well as dynamic elements, such as other vehicles and pedestrians. Sensor data, including cameras, LIDAR, and radar, are used to build this model. Through the use of the environmental model, the vehicle is capable of predicting the outcome of various actions. For instance, when a vehicle considers changing lanes, it uses its model to predict the behavior of surrounding vehicles to determine the safest and most efficient way to make the change. The model assists the vehicle in planning its route and making strategic decisions. To minimize travel time, avoid congestion, and enhance safety, it evaluates different routes and actions. Simulation allows the vehicle to select the best course of action by simulating various scenarios before implementation in the real world. The vehicle, for example, may use the model to simulate different actions in the event of a busy intersection, such as waiting for a gap in traffic or taking an alternate route. Considering the potential outcomes of each action, the vehicle can make an informed decision that balances efficiency with safety. In addition to improving the ability of autonomous vehicles to navigate safely and efficiently in real-world conditions, this model-based approach enables them to make complex decisions with a high level of accuracy. As a result of continuously refining the model based on new data, the vehicle is able to enhance its decision-making capabilities over time, thereby improving performance and enhancing safety on the road.
@@
There are several advantages to using model-based methods over methods that do not use models. By simulating future states and rewards, they can plan and evaluate different action sequences without interacting directly with the environment. It is believed that this capability may lead to a faster convergence to an optimal policy, since learning can be accelerated by leveraging the model’s predictions. A model-based approach can also adapt more quickly to changes in the environment, since it enables the model to be updated and re-planned accordingly. Although model-based methods have many advantages, they also face a number of challenges, primarily in regards to accuracy and computational cost. In order to create an accurate model of the environment, a high-fidelity model needs to be created. Moreover, the planning process may be computationally expensive, especially in environments with a large number of states and actions. However, advances in computing power and algorithms continue to improve the feasibility and performance of model-based methods, making them a valuable approach in RL [moerland2023model].

### v2

@@ 2.5 Policy Evaluation (Prediction)
Policy evaluation, also known as prediction, involves computing the state-value function $v_{\pi}$ for a given policy $\pi$. This process assesses the expected return when following policy $\pi$ from each state. The state-value function $v_{\pi}(s)$ is defined as the expected return starting from state $s$ and following policy $\pi$:
$v_{\pi}(s)\equiv\mathbb{E}_{\pi}[\mathcal{R}_{t+1}+\gamma\mathcal{G}_{t+1}\mid\mathcal{S}_{t}=s]$ (25)
@@
$v_{\pi}(s)=\mathbb{E}_{\pi}[\mathcal{R}_{t+1}+\gamma v_{\pi}(\mathcal{S}_{t+1})\mid\mathcal{S}_{t}=s]=\sum_{a}\pi(a\mid s)\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)[r+\gamma v_{\pi}(s^{\prime})]$ (26)
In these equations, $\pi(a\mid s)$ denotes the probability of taking action $a$ in state $s$ under policy $\pi$. The existence and uniqueness of $v_{\pi}$ are guaranteed if $\gamma<1$ or if all states eventually terminate under $\pi$. Dynamic Programming (DP) algorithm updates are termed ”expected updates” because they rely on the expectation over all potential next states, rather than just a sample [1].
@@ 2.6 Policy Improvement
The purpose of calculating the value function for a policy is to identify improved policies. Assuming $v_{\pi}$ for a deterministic policy $\pi$, for a state $s$, should we alter the policy to select action $a\neq\pi(s)$? We know the effectiveness of adhering to the existing policy from state $s$ ($v_{\pi}(s)$), but would transition to a new policy yield a superior outcome? We can answer this by selecting action $a$ in $s$ and then following $\pi$. To determine if a policy can be improved, we compare the value of taking a different action $a$ in state $s$ with the current policy. This is done using the action value function $q_{\pi}(s,a)$:
$q_{\pi}(s,a)\equiv\mathbb{E}[\mathcal{R}_{t+1}+\gamma v_{\pi}(\mathcal{S}_{t+1})\mid\mathcal{S}_{t}=s,\mathcal{A}_{t}=a]=\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)[r+\gamma v_{\pi}(s^{\prime})]$ (27)
The key criterion is whether this value exceeds $v_{\pi}(s)$. If $q_{\pi}(s,a)>v_{\pi}(s)$, consistently choosing action $a$ in $s$ is more advantageous than following $\pi$, leading to an improved policy $\pi^{\prime}$.
@@ 2.7 Policy Improvement Theorem
The policy improvement theorem states that if $q_{\pi}(s,\pi^{\prime}(s))\geq v_{\pi}(s)$ for all states $s$, then the new policy $\pi^{\prime}$ is at least as good as the original policy $\pi$. Formally, it is expressed as
@@
If $\pi^{\prime}$ achieves greater or equal expected return from all states $s\in\mathcal{S}$:
@@
If there is strict inequality at any state, $\pi^{\prime}$ is superior to $\pi$. Extending this to all states and actions, selecting the action that maximizes $q_{\pi}(s,a)$. The new policy $\pi^{\prime}$ is obtained by selecting the action that maximizes the action value function $q_{\pi}(s,a)$.
$\pi^{\prime}(s)\equiv\arg\max_{a}q_{\pi}(s,a)=\arg\max_{a}\mathbb{E}[\mathcal{R}_{t+1}+\gamma v_{\pi}(\mathcal{S}_{t+1})\mid\mathcal{S}_{t}=s,\mathcal{A}_{t}=a]=\arg\max_{a}\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)[r+\gamma v_{\pi}(s^{\prime})]$ (30)
Policy improvement creates a new policy that enhances an initial policy by adopting a greedy approach based on the value function. Assuming $\pi^{\prime}$ is equally effective as $\pi$ but not superior, then $v_{\pi}=v_{\pi^{\prime}}$ ensures that for all states $s\in\mathcal{S}$. The relationship between the optimal state value function $v_{*}(s)$ and the optimal action value function $q_{*}(s,a)$ is given by the following equation:
$v_{\pi^{\prime}}(s)=\max_{a}\mathbb{E}[\mathcal{R}_{t+1}+\gamma v_{\pi^{\prime}}(\mathcal{S}_{t+1})\mid\mathcal{S}_{t}=s,\mathcal{A}_{t}=a]=\max_{a}\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)[r+\gamma v_{\pi^{\prime}}(s^{\prime})]$ (31)
Policy improvement yields a superior policy unless the initial policy is already optimal. This concept extends to stochastic policies. Stochastic policies introduce a set of probabilities for actions, with the action most aligned with the greedy policy assigned the highest probability.
@@ 2.8 Policy Iteration
After enhancing a policy $\pi$ using $v_{\pi}$ to derive an improved policy $\pi^{\prime}$, compute $v_{\pi^{\prime}}$ and further refine it to obtain a superior policy $\pi^{\prime\prime}$. This process generates a sequence of improving policies and corresponding value functions. The process of policy iteration involves alternating between policy evaluation and policy improvement to obtain a sequence of improving policies and value functions:
@@
Each policy in this sequence is a marked improvement over its predecessor unless the preceding one is already optimal. Given a finite MDP, this iterative process converges to an optimal policy and value function in a finite number of iterations. This method is called Policy Iteration. Policy iteration involves two processes: policy evaluation aligns the value function with the current policy, and policy improvement makes the policy greedier based on the value function. These processes iteratively reinforce each other until an optimal policy is obtained.
2.9 Value Iteration
@@
One limitation of policy iteration is that each iteration requires policy evaluation, often necessitating multiple passes through the entire state set [12]. To address this, policy evaluation can be abbreviated without losing convergence guarantees. This method, known as value iteration, terminates policy evaluation after a single sweep. It combines policy improvement with a truncated form of policy evaluation. Value iteration merges one pass of policy evaluation with policy improvement in each iteration, ensuring convergence to an optimal policy for discounted finite MDPs [13]. The update rule for value iteration is given in Equation 33.
$v_{k+1}(s)\equiv\max_{a}\mathbb{E}[\mathcal{R}_{t+1}+\gamma v_{k}(\mathcal{S}_{t+1})\mid\mathcal{S}_{t}=s,\mathcal{A}_{t}=a]=\max_{a}\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)[r+\gamma v_{k}(s^{\prime})]$ (33)
@@
In value iteration, the key advantage is its efficiency, as it reduces the computational burden by merging policy evaluation and improvement into a single update step. This method is particularly useful for large state spaces where full policy evaluation at each step of policy iteration is computationally prohibitive [1]. Additionally, value iteration can be implemented using a synchronous update approach, where all state values are updated simultaneously, or an asynchronous update approach, where state values are updated one at a time, potentially allowing for faster convergence in practice. Another notable aspect of value iteration is its robustness to initial conditions. Starting from an arbitrary value function, value iteration iteratively refines the value estimates until convergence, making it a reliable method for finding optimal policies even when the initial policy is far from optimal [14]. Furthermore, value iteration provides a foundation for more advanced algorithms by illustrating the principle of bootstrapping, where the value of a state is updated based on the estimated values of successor states. This principle is central to many RL algorithms that seek to balance exploration and exploitation in dynamic and uncertain environments [15].
The concepts introduced in this section—ranging from bandits and MDPs to value functions and policies—provide the mathematical and conceptual tools required to understand RL methods. In the next section, we explore how these foundational ideas evolve into core RL algorithms, bridging theory and application.
3 Core RL Methods
@@
Understanding the various methodologies and concepts within RL is essential for the effective design and implementation of RL algorithms. Methods in RL can be classified as either off-policy or on-policy, and as model-free and model-based. These categories offer different approaches and techniques for learning from interactions with the environment.
3.1 Model-free & Model-based methods
@@
Model-free methods determine the optimal policy or value function directly without constructing a model of the environment. There is no requirement for them to know transition probabilities and rewards, as they learn entirely from observed states, actions, and rewards. Compared with model-based methods, model-free methods are simpler to implement, relying on experience-based learning. There are two primary types: Value-based and Policy-based methods. The former focus on learning the action-value function to derive an optimal policy. For instance, Q-learning (discussed in section 4) is an off-policy algorithm that learns the value of the optimal policy independently of the agent’s actions by using a max operator in its update rule. SARSA (also discussed in section 4), on the other hand, is an on-policy algorithm that updates its Q-values based on the actions actually taken by the policy. Both methods update their action-value estimates based on the Bellman equation until convergence. In contrast, policy-based methods, like REINFORCE (discussed in section 4), work by directly learning the policy without explicitly learning a value function. These methods adjust the policy parameters directly by following the gradient of the expected reward. This approach is particularly useful in environments with high-dimensional action spaces where value-based methods may not be effective. Policy-based methods are also capable of handling stochastic policies, providing a natural framework for dealing with uncertainty in action selection. In addition to these primary types, there are also hybrid approaches that combine value-based and policy-based methods, such as Actor-Critic algorithms (which will be discussed in section 4). These methods consist of two main components: an actor that updates the policy parameters in a direction suggested by the critic, and a critic that evaluates the action-value function. Combining both types of learning is intended to provide more stable and efficient learning [16].
@@
Another significant advancement in model-free methods is the development of Deep RL (DRL) By integrating deep neural networks with traditional RL algorithms, methods such as Deep Q-Networks (DQN) [17] and Proximal Policy Optimization (PPO) [18] have achieved remarkable success in complex, high-dimensional environments, including games and robotic control tasks. The advancement of these technologies has opened up new possibilities for the application of RL to real-world problems, enabling the demonstration of robust performance in domains which were previously intractable. It is beyond the scope of this paper to discuss these algorithms, and wee refer you to [19, 20, 21, 22] to understand DRL deeply and effectively.
@@
It is possible to predict the outcomes of actions using model-based methods, which facilitate strategic planning and decision-making. The use of these methods enhances learning efficiency by providing opportunities for virtual experimentation, despite the complexity of developing and refining accurate models [7]. Autonomous driving systems are an example of how model-based methods can be applied in the real world. As autonomous vehicles navigate in dynamic environments, obstacle avoidance, and optimal routing must be made in real time. Autonomous vehicles create detailed models of their environment. These models include static elements, such as roads and buildings, as well as dynamic elements, such as other vehicles and pedestrians. Sensor data, including cameras, LIDAR, and radar, are used to build this model. Through the use of the environmental model, the vehicle is capable of predicting the outcome of various actions. For instance, when a vehicle considers changing lanes, it uses its model to predict the behavior of surrounding vehicles to determine the safest and most efficient way to make the change. The model assists the vehicle in planning its route and making strategic decisions. To minimize travel time, avoid congestion, and enhance safety, it evaluates different routes and actions. Simulation allows the vehicle to select the best course of action by simulating various scenarios before implementation in the real world. The vehicle, for example, may use the model to simulate different actions in the event of a busy intersection, such as waiting for a gap in traffic or taking an alternate route. Considering the potential outcomes of each action, the vehicle can make an informed decision that balances efficiency with safety. In addition to improving the ability of autonomous vehicles to navigate safely and efficiently in real-world conditions, this model-based approach enables them to make complex decisions with a high level of accuracy. As a result of continuously refining the model based on new data, the vehicle is able to enhance its decision-making capabilities over time, thereby improving performance and enhancing safety on the road.
@@
There are several advantages to using model-based methods over methods that do not use models. By simulating future states and rewards, they can plan and evaluate different action sequences without interacting directly with the environment. It is believed that this capability may lead to a faster convergence to an optimal policy, since learning can be accelerated by leveraging the model’s predictions. A model-based approach can also adapt more quickly to changes in the environment, since it enables the model to be updated and re-planned accordingly. Although model-based methods have many advantages, they also face a number of challenges, primarily in regards to accuracy and computational cost. In order to create an accurate model of the environment, a high-fidelity model needs to be created. Moreover, the planning process may be computationally expensive, especially in environments with a large number of states and actions. However, advances in computing power and algorithms continue to improve the feasibility and performance of model-based methods, making them a valuable approach in RL [23].